Hello,i still have another age problem,I hope you help me!Thank you!
:Ann is 5 years older than Kristine.In 10 years,Kristine's age is 24 more than the difference of their ages will be.How old are they now?


Answer by rothauserc(4718)

let x be Ann's age and y be Kristine's age, then
x = y +5
y+10 = (x+10 - y+10)+24
substitute for x in second equation
y+10 = (y+5+10 -y-10) + 24
y+10 = 5 +24
y+10 = 29
y = 19
x = 24
Ann's age is 24 and Kristine's age is 19